# Data assimilation for networks of coupled oscillators: Inferring unknown model parameters from partial observations

### v2

@@ Abstract
Inferring the state and unknown parameters of a network of coupled oscillators, such as neurons in the brain, is of utmost importance. This task is made harder when only partial and noisy observations are available, which is a typical scenario in realistic high-dimensional systems. The general task of inference falls under data assimilation, and a commonly used assimilation method is the Ensemble Kalman Filter. Employing network-specific localization of the forecast covariance, an Ensemble Kalman Filter with state space augmentation is shown to yield highly accurate estimates of both the oscillator phases and unknown model parameters in the case where only a subset of oscillator phases are observed. In contrast, standard data assimilation methods yield poor results. We demonstrate the effectiveness of our approach for Kuramoto oscillators and for networks of theta neurons, using a variety of network topologies.
@@ 1 Introduction
Many natural phenomena and engineering applications can be described as networks of coupled oscillators, for example, the firing of neurons in the brain [1, 2, 3] and the dynamics of power grids [4, 5, 6, 7]. However, mathematical models for these systems are often incomplete, with unknown parameters and simplifying physical assumptions. Scientists are now seeking data-driven methods to improve estimates of the dynamic state of these systems using observational data, and to simultaneously estimate unknown parameters.
@@
A pertinent question is: Can one accurately estimate all the oscillators’ phases as well as unknown model parameters if only a subset of the oscillator phases are observed? For example, in a power grid, we may be able to observe all the power stations, but none of the consumers. The framework of data assimilation (DA) provides a unifying framework that estimates a system’s state and unknown parameters by optimally combining two uncertain pieces of information – the output of our imperfect model (the model forecast) and noisy partial observations of the system at discrete time intervals.
@@
We focus specifically on the Ensemble Kalman Filter (EnKF) [8, 9, 10, 11, 12, 13, 14] as a DA method. Although the framework of Kalman filters can strictly only be applied to linear systems, the EnKF has been successfully used in numerous non-linear settings [15], in particular, weather forecasting [8, 9, 10, 11, 12, 13, 14]. To date, there have been only a few applications of DA to the broad class of non-linear systems that model coupled oscillators on complex networks [16, 17]. These applications, however, assume that all nodes are observed. Indeed, we find that the standard EnKF performs poorly for networks of coupled oscillators unless all nodes are observed or very large ensembles are used. Localization has been recognized as an essential component for successful performance of EnKFs [18, 19]. In order to employ localization, a notion of distance is required. In applications involving partial differential equations on spatial domains, for example in weather forecasting, the Euclidean distance is appropriate and localization has shown to be highly efficient in mitigating detrimental small ensemble size effects. It is less clear how best to incorporate “distance” and localization for dynamics on networks. We introduce a novel localization method, specific to dynamics on networks, which renders the EnKF highly effective at estimating both oscillator phases and unknown model parameters, even when a small ensemble is used and only a fraction of the phases are observed. This sets DA apart from other data-driven parameter estimation methods that require all oscillator phases to be observed [20, 21, 22, 23]. We demonstrate our method for the Kuramoto model [24, 25, 26, 27, 28, 29, 30, 31] which captures the synchronization phenomena observed in many coupled oscillator systems, and for networks of theta neurons [32, 33, 34, 35] which model neural dynamics of the brain.
@@ 3 Ensemble Kalman Filter (EnKF) for coupled oscillators on networks
To solve the inference problem of estimating the state of a system and unknown parameters of the underlying model in the case that only some of the state variables are observed we employ the Ensemble Kalman Filter (EnKF). Kalman filters estimate the posterior distribution of the state given a model forecast and noisy observations. At each discrete time-step $t_{n}$, the EnKF optimally combines a forecast ensemble $X_{n}^{\rm{f}}$ obtained by independently simulating the full non-linear model from a previous DA step, and noisy observations $Y_{n}$ to yield an improved ‘analysis’ ensemble.
@@
In a Kalman filter, a Gaussian approximation of the posterior distribution is performed and one seeks expressions for the mean $X^{\rm a}$ and the covariance matrix $P^{\rm a}$ of the posterior. We present here the algorithmic details to compute $X^{\rm a}$ and $P^{\rm a}$. For more details and derivations the reader is referred to [8, 9, 10, 11, 12, 13, 14]. The analysis $X^{\rm a}_{n}$ at time $t_{n}$ is the optimal estimate of the state given a forecast $X^{\rm f}_{n}$ and an observation $Y_{n}$. Treating $X^{\rm f}_{n}$ and $X^{\rm a}_{n}$, $n\geq 0$ as random variables and assuming a Gaussian distribution for $X_{n+1}^{\rm f}$, the analysis step for the mean $\overline{X}^{\rm a}_{n}$ is given by
@@
In finite ensembles members may align, on short time scales, with the most unstable direction leading to an ensemble collapse in which many ensemble members become indistinguishable. This leads to filter divergence in which the filter believes its own forecast since $P^{\rm f}$ is small, and the forecast is not corrected by new incoming observations. To mitigate against such finite ensemble size effects we employ, as is typically done, covariance inflation with $P^{\rm f}_{n}\to\delta P^{\rm f}_{n}$ with $\delta=1.001$ [41]. This multiplicative inflation does not effect the ensemble mean but increases the forecast error covariance.
@@ 3.1 Splitting the analysis step into phase and parameter components
3.2 Spurious correlations and localization
A source of error in ensemble-based methods arises due to spurious correlations caused by the finite size of the ensemble. If two variables are not correlated, the Monte-Carlo approximation of the covariance yields entries of ${\mathcal{O}(1/\sqrt{M}})$ compared to the true entry which is 00. One could reduce the effect of spurious correlations by using a very large ensemble size $M$, but that is computationally prohibitive for high-dimensional systems. An alternative approach is so-called “localization”, whereby topological information is used to predetermine and suppress spurious correlations [18], which allows for a smaller ensemble to be used. For example, in weather forecasting, correlations are deemed spurious if the grid-points are geographically distant. It is less clear how to best incorporate localization for dynamics on networks.
@@ 3.2 Spurious correlations and localization
where $\circ$ denotes the Schur (elementwise) product, and $\mathcal{L}$ is the localization matrix. Standard localization functions, such as Gaspari-Cohn localization [18, 42], are not suitable for networks, because they rely on Euclidean geometries. We discuss this in detail in Section 3.4. Instead we require $\mathcal{L}$ to have the following properties:
  • $\mathcal{L}$ encodes the network connectivity, i.e., $\mathcal{L}_{ij}$ is large if nodes $i$ and $j$ are well connected, and small otherwise,
  • $\mathcal{L}$ is a correlation matrix, i.e., positive semi-definite and $\mathcal{L}_{ii}=1$.
Consider the (symmetric positive semi-definite matrix) covariance matrix, $\mathcal{E}=\exp(\lambda A)$, where $\lambda$ is a tunable parameter and $A$ is the network adjacency matrix (assumed to be symmetric). The matrix
@@
is a weighted sum, with decreasing weights, of powers $A^{n}$. The entries $(A^{n})_{ij}$ count the number of paths of length $n$ between nodes $i$ and $j$. Hence, $\mathcal{E}$ encodes the network topology. A correlation matrix $L$ is obtained by normalizing $\mathcal{E}$ according to
@@
Figure 3: (a) Average correlation matrix $Q$ for a large ensemble ($M=10100$), and (b) for a small ensemble ($M=101$), obtained by applying the standard EnKF to the Kuramoto model (1). The augmented state space is $(\phi_{1,\dots,50},\omega_{1,\dots,50})$, hence the block matrix substructure (indicated by red lines) in $Q$ with $\phi$-$\phi$ correlations, $\phi$-$\omega$ correlations, $\omega$-$\phi$ correlations and $\omega$-$\omega$ correlations. The Kuramoto model parameters are as in Fig. 1. (c) The localization matrix $\mathcal{L}$ (23) for the associated ring topology using $\lambda=0.46$ as determined in Section 3.3.
@@ 3.3 Choosing the localization parameter $\lambda$
The parameter $\lambda$ in the localization matrix $L$ (21), (22) is important for controlling the rate of decay of correlations for distant nodes. The optimal value of $\lambda$ could be found through numerical optimization, minimizing the RMS error of the DA algorithm across many simulations. Here we present a cheap heuristic method to choose a value of $\lambda$ at first for ring network topologies, and then we will generalize the method to random network topologies. Our heuristic is based on the typical decay of correlations that is observed from DA for large ensembles.
@@ 3.3.1 Ring network topologies
Considering a ring topology with connectivity radius $r$, we find that the correlations for large ensembles are negligible beyond a radius of $2r$. That is, for a node $j$, only the nodes $k$ that have shortest path length from $j$ to $k$ less than or equal to $2$ have significant correlations, and all other correlations can be considered spurious. This is shown in Fig. 4, which shows the average correlation as a function of distance from the node, scaled by the connectivity radius $r$. This data is obtained from the standard EnKF applied to the Kuramoto model with $N=50$ and with a large ensemble $M=100(2N+1)$, as in Fig. 3(a) which shows correlations for $r=3$. Fig. 4 shows that correlations become spurious beyond a distance of $2r$ ($d/r=2$). Using this information, considering node $j=1$, the nodes within a radius of $2r$, and which will have significant correlations are $k=1,\dots,2r+1$ and $k=N-r+1,\dots,N$, and all other nodes, i.e., $k=2r+2,\dots,N-r$ will have spurious correlations. Knowing which correlations we wish to suppress, the next question is how much do we need to suppress them? For this, we choose an $10<\epsilon<$ 1, which will be the upper bound for the entries in $L$ (22) beyond a radius of $2r$ from each node. Here we have chosen $\epsilon=0.1$. Due to the symmetry of the ring topology, and the fact that $\exp(\lambda A)$ decays away from its diagonal, we require that
@@
where $\mathcal{E}=\exp(\lambda A)$ as in (21). Changing the inequality in (24) to an equality gives an equation that can be solved (numerically) for $\lambda$ in terms of $r$ and $\epsilon$. This is demonstrated schematically in Fig. 5, where the orange squares are the first spurious correlations, with $L_{ij}=\epsilon$. In particular, in the first row we have $L_{1,2r+2}=\epsilon$.
Figure 4: Average forecast correlation (averaging $P_{k}^{\rm f}$ over all nodes and all time steps) as a function of distance $d$ from the node, scaled by the coupling radius $r$, for $r=2,3,5,10$. All results are obtained from runs of the standard EnKF applied to the Kuramoto model ($N=50$, $\kappa=80/r$, $\omega_{i}\sim\mathcal{N}(0,0.1)$ with a ring topology), with a large ensemble $M=100(2N+1)$. Note that a coupling strength $\kappa=80/r$ is used to maintain a constant effective coupling strength $\kappa/\langle k\rangle$, where $\langle k\rangle$ is the mean degree.
Figure 5: Schematic diagram for the condition (24) that defines $\lambda$. Shown is the top right corner of the matrix $L$, the gray squares indicate the underlying adjacency matrix (here a ring with $r=2$), the green squares indicate correlations that should not be considered spurious because they are within a distance of $2r$ from the node, and orange/red squares indicate correlations that should be considered spurious because the distance from the node is greater than $2r$. The orange squares are those whose value will be $\epsilon$ for $\lambda$ satisfying (24) with equality, e.g., in the first row $L_{1,2r+2}=\epsilon$.

### v3

@@ Abstract
Inferring the state and unknown parameters of a network of coupled oscillators is of utmost importance. This task is made harder when only partial and noisy observations are available, which is a typical scenario in realistic high-dimensional systems. The general task of inference falls under data assimilation, and a commonly used assimilation method is the Ensemble Kalman Filter. Employing network-specific localization of the forecast covariance, an Ensemble Kalman Filter with state space augmentation is shown to yield highly accurate estimates of both the oscillator phases and unknown model parameters in the case where only a subset of oscillator phases are observed. In contrast, standard data assimilation methods yield poor results. We demonstrate the effectiveness of our approach for Kuramoto oscillators and for networks of theta neurons, using a variety of network topologies.
This is the accepted version to appear in Proceedings of the Royal Society A. The published version will be available at doi: 10.1098/rspa.2024.0813
@@ 1 Introduction
Many natural phenomena and engineering applications can be described as networks of coupled oscillators, for example, the firing of neurons in the brain [1, 2, 3, 4] and the dynamics of power grids [5, 6, 7, 8, 9]. However, mathematical models for these systems are often incomplete, with unknown parameters and simplifying physical assumptions. Scientists are now seeking data-driven methods to improve estimates of the dynamic state of these systems using observational data, and to simultaneously estimate unknown parameters.
@@
A pertinent question is: Can one accurately estimate all the oscillators’ phases as well as unknown model parameters if only a subset of the oscillator phases are observed? For example, in a power grid, we may be able to observe all the power stations, but none of the consumers. The framework of data assimilation (DA) provides a unifying framework that estimates a system’s state and unknown parameters by combining two uncertain pieces of information – the output of our imperfect model (the model forecast) and noisy partial observations of the system at discrete time intervals.
@@
We focus specifically on the Ensemble Kalman Filter (EnKF) [10, 11, 12, 13, 14, 15, 16] as a DA method. Although the framework of Kalman filters can strictly only be applied to linear systems [17], the EnKF has been successfully used in numerous non-linear settings [18], in particular, weather forecasting [10, 11, 12, 13, 14, 15, 16]. To date, there have been only a few applications of DA to the broad class of non-linear systems that model coupled oscillators on complex networks [19, 20]. These applications, however, assume that all nodes are observed. Indeed, we find that the standard EnKF performs poorly for networks of coupled oscillators unless all nodes are observed or very large ensembles are used. Covariance localization, whereby potentially spurious entries of the covariance matrix are set to zero, has been recognized as an essential component for successful performance of EnKFs [21, 22]. In order to employ localization, one needs to identify those entries of the sample covariance matrix which are overwhelmed by finite-sample noise [11, 23, 24, 25, 26]. If two variables are not correlated or are very weakly correlated, the corresponding entry of the sample covariance matrix should be small, but is easily contaminated by finite-sample noise. One often employs a notion of distance to identify which variables are weakly correlated: the larger the distance the less their mutual effect on each other. In applications involving partial differential equations on spatial domains, for example in weather forecasting, the Euclidean distance is appropriate and localization has shown to be highly efficient in mitigating detrimental small ensemble size effects. It is less clear how best to incorporate “distance” and localization for dynamics on networks. We introduce a novel localization method, specific to dynamics on networks, which renders the EnKF highly effective at estimating both oscillator phases and unknown model parameters, even when a small ensemble is used and only a fraction of the phases are observed. This sets DA apart from other data-driven parameter estimation methods that require all oscillator phases to be observed [27, 28, 29, 30]. We demonstrate our method for the Kuramoto model [31, 32, 33, 34, 35, 36, 37, 38] which captures the synchronization phenomena observed in many coupled oscillator systems, and for networks of theta neurons [39, 40, 41, 42] which model neural dynamics of the brain.
@@ 2.1 The Kuramoto model
The Kuramoto model (1) can be thought of as an inertia-less simplification of the dynamics of power grids, wherein the oscillators of the network are the power stations, consumers and buses, and power lines form the network edges. The dynamics of power grids is described by the swing equation which can be written as the second order Kuramoto model [7, 8, 9] which includes inertia. Real-time observational data can be obtained from the nodes in the power grid, such as their production or consumption, their operating frequency, and their phase difference from the local net phase [5, 6].
@@ 2.2 Networks of theta neurons
The true dynamics of the brain can be described by very high dimensional systems of delay differential equations, with a node for each of the billions of neurons. The dimensionality and complexity can be reduced by grouping sets of neurons into distinct regions of the brain, as well as simplifying the dynamics by ignoring, for instance, the delay. The theta neuron model (2) is particularly simple and not very realistic, but we remark that slightly more complicated models, such as the FitzHugh-Nagumo model, are able to accurately replicate brain dynamic phenomena such as the temporary synchronization that occurs during epileptic seizures [4]. For brain dynamics, electroencephalography (EEG) can yield real-time data about the activity within different regions of the brain, and from that, the local phase and local frequency of the regions can be determined [4, 45].
@@ 3 Ensemble Kalman Filter (EnKF) for coupled oscillators on networks
To solve the inference problem of estimating the state of a system and unknown parameters of the underlying model in the case that only some of the state variables are observed we employ the Ensemble Kalman Filter (EnKF). Kalman filters estimate the posterior distribution of the state given a model forecast and noisy observations. At each discrete time-step $t_{n}$, the EnKF combines a forecast ensemble $X_{n}^{\rm{f}}$ obtained by independently simulating the full non-linear model from a previous DA step, and noisy observations $Y_{n}$ to yield an improved ‘analysis’ ensemble.
@@
In a Kalman filter, a Gaussian approximation of the posterior distribution is performed and one seeks expressions for the mean $X^{\rm a}$ and the covariance matrix $P^{\rm a}$ of the posterior. We present here the algorithmic details to compute $X^{\rm a}$ and $P^{\rm a}$. For more details and derivations the reader is referred to [10, 11, 12, 13, 14, 15, 16]. The analysis $X^{\rm a}_{n}$ at time $t_{n}$ is an estimate of the state that combines a forecast $X^{\rm f}_{n}$ and an observation $Y_{n}$. Treating $X^{\rm f}_{n}$ and $X^{\rm a}_{n}$, $n\geq 0$ as random variables and assuming a Gaussian distribution for $X_{n+1}^{\rm f}$, the analysis step for the mean $\overline{X}^{\rm a}_{n}$ is given by
@@
Finite ensembles are prone to an underestimation of error covariances which may cause filter divergence in which the filter believes its own forecast since $P^{\rm f}$ is small, and the forecast is not corrected by new incoming observations. To mitigate against such finite ensemble size effects we employ, as is typically done, covariance inflation with $P^{\rm f}_{n}\to\delta P^{\rm f}_{n}$ with $\delta=1.001$ [49]. This multiplicative inflation does not affect the ensemble mean but increases the forecast error covariance. The exact causes of how finite ensemble sizes lead to underestimation of covariances is complex [11, 50, 51, 52]. In chaotic dynamical systems the underestimation of error covariances may be further exaggerated by ensemble members aligning, on short time scales, with the most unstable direction [53].
@@ 3.1 Splitting the analysis step into phase and parameter components
3.2 Network specific localization
A source of error in ensemble-based methods arises due to spurious correlations caused by the finite size of the ensemble. If two variables are not correlated, the Monte-Carlo approximation of the covariance yields entries of ${\mathcal{O}(1/\sqrt{M}})$ compared to the true entry which is 00. One could reduce the effect of spurious correlations by using a very large ensemble size $M$, but that is computationally prohibitive for high-dimensional systems. An alternative approach is so-called “localization”, whereby topological information is used to predetermine and suppress spurious correlations [21], which allows for a smaller ensemble to be used. For example, in weather forecasting, correlations between geographically distant grid-points should be very small and any large correlations are deemed spurious. It is less clear how to best incorporate localization for dynamics on networks.
@@ 3.2 Network specific localization
where $\circ$ denotes the Schur (elementwise) product, and $\mathcal{L}$ is the localization matrix. Standard localization functions, such as Gaspari-Cohn localization [21, 54], are not suitable for networks, because they cannot be embedded in $\mathbb{R}^{n}$ for any $n$. We discuss this in detail in Section 3.4. Instead we require $\mathcal{L}$ to have the following properties:
  • $\mathcal{L}$ encodes the network connectivity, i.e., $\mathcal{L}_{ij}$ is large if nodes $i$ and $j$ are well connected in the sense that they are direct neighbors and/or there are many paths of short length connecting them, and $\mathcal{L}_{ij}$ is small if nodes $i$ and $j$ are poorly connected,
  • $\mathcal{L}$ is a correlation matrix, i.e., symmetric, positive semi-definite and $\mathcal{L}_{ii}=1$.
As a natural choice of a matrix that has the two above properties, we propose the matrix exponential of the network adjacency matrix. Consider the (symmetric positive semi-definite matrix) covariance matrix, $\mathcal{E}=\exp(\lambda A)$, where $\lambda$ is a tunable parameter and $A$ is the network adjacency matrix (assumed to be symmetric). The matrix
@@
is a weighted sum, with decreasing weights, of powers $A^{n}$. The entries $(A^{n})_{ij}$ count the number of paths of length $n$ between nodes $i$ and $j$. Hence, $\mathcal{E}$ encodes the network topology. The entry $\mathcal{E}_{ij}$ quantifies the connectedness between nodes $i$ and $j$; $\mathcal{E}_{ij}$ is large if nodes $i$ and $j$ are directly connected, or connected via many paths with short length, and $\mathcal{E}_{ij}$ is close to zero otherwise. A correlation matrix $L$ is obtained by normalizing $\mathcal{E}$ according to
@@
Figure 3: (a) Time-averaged correlation matrix $Q$ for a large ensemble ($M=10100$), and (b) for a small ensemble ($M=101$), obtained by applying the standard EnKF to the Kuramoto model (1). The time-averaging is performed for a single realization. The augmented state space is $(\phi_{1,\dots,50},\omega_{1,\dots,50})$, hence the block matrix substructure (indicated by red lines) in $Q$ with $\phi$-$\phi$ correlations, $\phi$-$\omega$ correlations, $\omega$-$\phi$ correlations and $\omega$-$\omega$ correlations. Note that the absolute values of the correlations are shown on a logarithmic scale to more clearly illustrate the differences in magnitude between the smaller correlations. The Kuramoto model parameters are as in Fig. 1. (c) The localization matrix $\mathcal{L}$ (23) for the associated ring topology using $\lambda=0.46$ as determined in Section 3.3.
@@
We remark that there exist other localization matrices that satisfy the two desired properties above, and we discuss some alternative localization matrices in Section 3.4.
@@ 3.3 Choosing the localization parameter $\lambda$
The parameter $\lambda$ in the localization matrix $L$ (21), (22) is important for controlling the rate of decay of correlations for distant nodes. The optimal value of $\lambda$ could be found through numerical optimization, minimizing the RMS error of the DA algorithm across many simulations. Here we present a cheap heuristic method to choose a value of $\lambda$ at first for ring network topologies, and then we will generalize the method to random network topologies.
We note that the distance at which correlations become spurious, and, hence, the optimal choice of $\lambda$, will depend on the size of the ensemble $M$, such that larger ensembles are able to better estimate the weaker long range correlations [11]. The heuristic method that we present here is based on the typical decay of correlations as observed from DA for a large ensemble, which allows for a reliable estimation of the typical distance at which correlations become weak.
@@ 3.3.1 Ring network topologies
Considering a ring topology with connectivity radius $r$, we find that the correlations for large ensembles are weak beyond a radius of $2r$. That is, for a node $j$, only the nodes $k$ that have shortest path length from $j$ to $k$ less than or equal to $2$ have significant correlations. This is shown in Fig. 4, which shows the average correlation as a function of distance from the node, scaled by the connectivity radius $r$. This data is obtained from the standard EnKF applied to the Kuramoto model with $N=50$ and with a large ensemble $M=100(2N+1)$, as in Fig. 3(a) which shows correlations for $r=3$. Fig. 4 shows that correlations become weak beyond a distance of $2r$ ($d/r=2$). The rapid decay of correlation that we have observed is expected due to the pairwise coupling of nodes in both the Kuramoto model (1) and theta neuron model (2), which restricts the flow of information. Using this information, considering node $j=1$, the nodes within a radius of $2r$, and which will have significant correlations are $k=1,\dots,2r+1$ and $k=N-r+1,\dots,N$, and all other nodes, i.e., $k=2r+2,\dots,N-r$ should have weak correlations. However, these nodes that should have weak correlations may in fact have large, spurious, correlations due to a small ensemble size $M$. These potentially spurious correlations are those which should be suppressed. Knowing which correlations we wish to suppress, the next question is how much do we need to suppress them? For this, we choose an $10<\epsilon<$ 1, which will be the upper bound for the entries in $L$ (22) beyond a radius of $2r$ from each node. Here we have chosen $\epsilon=0.1$. Due to the symmetry of the ring topology, and the fact that $\exp(\lambda A)$ decays away from its diagonal, we require that
@@
where $\mathcal{E}=\exp(\lambda A)$ as in (21). Changing the inequality in (24) to an equality gives an equation that can be solved (numerically) for $\lambda$ in terms of $r$ and $\epsilon$. This is demonstrated schematically in Fig. 5, where the orange squares are the first weak correlations, with $L_{ij}=\epsilon$. In particular, in the first row we have $L_{1,2r+2}=\epsilon$.
Figure 4: Average forecast correlation (averaging $P_{k}^{\rm f}$ over all nodes and all time steps) as a function of distance $d$ from the node, scaled by the coupling radius $r$, for $r=2,3,5,10$. All results are obtained from runs of the standard EnKF applied to the Kuramoto model ($N=50$, $\kappa=80/r$, $\omega_{i}\sim\mathcal{N}(0,0.1)$ with a ring topology), with a large ensemble $M=100(2N+1)$. Note that a coupling strength $\kappa=80/r$ is used to maintain a constant effective coupling strength $\kappa/\langle k\rangle$, where $\langle k\rangle$ is the mean degree, which ensures that the dynamics of the Kuramoto model are equivalent for all $r$.
Figure 5: Schematic diagram for the condition (24) that defines $\lambda$. Shown is the top right corner of the matrix $L$, the gray squares indicate the underlying adjacency matrix (here a ring with $r=2$), the green squares indicate correlations that should be considered significant because they are within a distance of $2r$ from the node, and orange/red squares indicate correlations that should be considered weak and potentially spurious because the distance from the node is greater than $2r$. The orange squares are those whose value will be $\epsilon$ for $\lambda$ satisfying (24) with equality, e.g., in the first row $L_{1,2r+2}=\epsilon$.